don't know if you can help but here is the question?
Suppose that 4e^(5x) + 8 = 7e^(5x) - 1. Solve for x

Answer by TP(29)

7e^(5x) - 1 = 4e^(5x) + 8
Add 1 to each side
7e^(5x) = 4e^(5x) + 9
Subtract 4e^(5x) from both sides
3e^(5x) = 9
Divide both sides by 3
e^(5x) = 3
Take logs of both sides
In[e^(5x)] = In(3)
So 5xIn(e) = In(3)
So 5x = In(3) (since In(e) = 1)
Divide both sides by 5
Hence x = [In(3)]/5 = [1.098612289]/5 = 0.219722457 ( to 9 dec. places) ANS